OpenStudy (anonymous):

What is the simplest form of the radical expression? 4^sqrt 625x^8y^12

OpenStudy (jdoe0001):

\(\large \bf \sqrt[4]{625x^8y^{12}}\\ \quad \\ \textit{keep in mind that }625=5^4\qquad x^8=x^4\cdot x^4\qquad y^{12}=y^4\cdot y^4\cdot y^4\)

OpenStudy (jdoe0001):

\(\bf \large \sqrt[4]{625x^8y^{12}}\\ \quad \\ \textit{keep in mind that }625=5^4\qquad x^8=x^4\cdot x^4\qquad y^{12}=y^4\cdot y^4\cdot y^4\\ \quad \\ \sqrt[4]{625x^8y^{12}}\implies \large \sqrt[{\color{red}{ 4}}]{5^4x^4x^4y^4y^4y^4}\)

OpenStudy (anonymous):

That doesn't fit any of my choices. My choices are A. \[5x^2\left| y^3 \right|\] B. \[25\left| x^2 \right|y^3\] C. \[5x^4\left| y^3 \right|\] D. \[25\left| x^4y^3 \right|\]

OpenStudy (anonymous):

A

OpenStudy (whpalmer4):

Yes, because we don't know the values of x and y, we can't do the simplification of y^12 quite so straightforwardly because y^3 will have a different value if y is negative.

OpenStudy (anonymous):

Thank you!

OpenStudy (whpalmer4):

For example, if we use y=1, x=1 for ease of arithmetic, the original expression works out to be 5. Same for y=-1, x=1. If we simplify it to 5x^2y^3 as one might do when unaware of this wrinkle, with y=1,x=1 we still get 5, but with y=-1,x=1, we get -5! Our simplification would be incorrect if \(y<0\). That's why writing it as \(5x^2|y^3|\) is done.